OpenStudy (anonymous):

Given two spheres, if the volume of the first sphere is 4/3 pie, and the volume of the second sphere is 36 pie, which is the relation of the second radius to the first? A.Squared B.Tripled C.Double

OpenStudy (anonymous):

The answer is B because The volume of the sphere is V = (4/3)(pi)(r^3) It is easy to see the radius of the small sphere is r = 1 The radius of the large sphere is V = (4/3)(pi)(r^3) = 36 pi r^3 = 36 (3/4) = 27 r = 3 We also can use the theorem Lengths Areas Volumes a/b (a/b)^2 (a/b)^3 However, in this specific problem, solving directly is easier.

OpenStudy (anonymous):

Thank you so much!! :)) Gaine A Fan!

OpenStudy (anonymous):

*Gained

OpenStudy (anonymous):

:)